# Robust First and Second-Order Differentiation for Regularized Optimal Transport

### v1

@@ 4 Spectral analysis of the H-matrix
This section analyzes the spectrum of the matrice $H(\Pi)$ in (2), which plays a crucial role in the computation of the derivative and the Hessian. Recall that $\Pi\in\mathbb{R}^{M\times N}$ is a coupling matrix if its entries are nonnegative and $\boldsymbol{\mu}=\Pi\mathbbm{1}_{N}$ and $\boldsymbol{\nu}=\Pi^{\top}\mathbbm{1}_{M}$ are marginal probability vectors. We say that it is a positive coupling matrix if its entries are all positive.
We consider two common types of coupling matrices $\Pi$: (i) positive coupling matrices, since the coupling matrix $\Pi^{*}$ obtained from the Sinkhorn algorithm in practice is always positive due to the entropy regularization, [13, 24]; and (ii) coupling matrix with uniform marginal distributions, i.e., $\boldsymbol{\mu}=\mathbbm{1}_{M}/M$ and $\boldsymbol{\nu}=\mathbbm{1}_{N}/N$, since they arise in most applications with randomly sampled data.
We show first that $H(\Pi)$ is singular with a simple zero eigenvalue for any positive coupling matrix. For coupling matrices with uniform marginal distributions (including permutation matrices), we analytically calculate the eigenvalues of $H(\Pi)$ in Section 4.2. In particular, we establish lower and upper bounds for the condition number of $H$-matrix in terms of the spectral gap of $\Pi$ in Theorem 9. Section 4.3 extends the bounds for the condition number to the H-matrix computed by the Sinkhorn algorithm of EOT distance with early-stopping.
@@ 4.1 General positive coupling matrices
We show first that the H-matrix $H(\Pi)$ is singular with a simple zero eigenvalue for any positive coupling matrix $\Pi$. As a result, we call $\kappa(H)=\frac{\lambda_{1}(H)}{\lambda_{N+M-1}(H)}$, the condition number of the matrix $H=H(\Pi)$.
Theorem 5 (Simple zero eigenvalue for the $H$-matrix).
For any positive coupling matrix $\Pi\in\mathbb{R}^{M\times N}_{>0}$, the smallest eigenvalue of $H(\Pi)$ is zero and it is simple, with eigenvector $\mathbf{q}_{0}=\left[\begin{array}[]{ccccccccc}\mathbbm{1}_{M}\\ -\mathbbm{1}_{N}\end{array}\right]$.
Its proof relies on the next Perron-Frobenius type lemma, which shows that the largest eigenvalue of the matrix $\text{diag}(\boldsymbol{\nu})^{-1}\Pi^{\top}\,\text{diag}(\boldsymbol{\mu})^{-1}\Pi$ is $1$ and is simple.
@@
For any coupling matrix with strictly positive entries $\Pi\in\mathbb{R}^{M\times N}_{>0}$, the largest eigenvalue of matrix $\text{diag}(\boldsymbol{\nu})^{-1}\Pi^{\top}\,\text{diag}(\boldsymbol{\mu})^{-1}\Pi$ is $\lambda=1$ and has multiplicity one, with eigenvector $\mathbbm{1}_{N}$. Similarly, the largest eigenvalue of the matrix $\text{diag}(\boldsymbol{\mu})^{-1}\Pi\,\text{diag}(\boldsymbol{\nu})^{-1}\Pi^{\top}$ is $\lambda=1$ and has multiplicity one, with eigenvector $\mathbbm{1}_{M}$.
@@
It is clear that 00 is an eigenvalue with eigenvector $\mathbf{q}_{0}$. We show next that any other eigenvector of 00 must be $\mathbf{q}_{0}$ up to a scalar factor. Note that the vector $\left[\begin{array}[]{ccccccccc}\boldsymbol{w}\\ \boldsymbol{v}\end{array}\right]$ is the eigenvector of 00 if and only if
$\left[\begin{array}[]{ccccccccc}\text{diag}(\boldsymbol{\mu})&\Pi\\ \Pi^{\top}&\text{diag}(\boldsymbol{\nu})\end{array}\right]\left[\begin{array}[]{ccccccccc}\boldsymbol{w}\\ \boldsymbol{v}\end{array}\right]=\mathbb{0}_{M+N}\Leftrightarrow\left[\begin{array}[]{ccccccccc}\text{diag}(\boldsymbol{\mu})\boldsymbol{w}+\Pi\boldsymbol{v}=0\\ \text{diag}(\boldsymbol{\nu})\boldsymbol{v}+\Pi^{\top}\boldsymbol{w}=0\end{array}\right],$ (45)
@@
$\boldsymbol{v}$ $=-\text{diag}(\boldsymbol{\nu})^{-1}\Pi^{\top}\boldsymbol{w}=\text{diag}(\boldsymbol{\nu})^{-1}\Pi^{\top}\text{diag}(\boldsymbol{\mu})^{-1}\Pi\boldsymbol{v},$
$\boldsymbol{w}$ $=-\text{diag}(\boldsymbol{\mu})^{-1}\Pi\boldsymbol{v}=\text{diag}(\boldsymbol{\mu})^{-1}\Pi\text{diag}(\boldsymbol{\nu})^{-1}\Pi^{\top}\boldsymbol{w}.$
By Lemma 6, $\boldsymbol{v}=a\mathbbm{1}_{N}$ and $\boldsymbol{w}=b\mathbbm{1}_{M}$ for some nonzero constant $a,b$ for each equation to hold. Plugging back to (45), we have $a=-b$, and $\left[\begin{array}[]{ccccccccc}\boldsymbol{w}\\ \boldsymbol{v}\end{array}\right]=a\mathbf{q}_{0}$. Thus, the zero eigenvalue of $H$ is simple. ∎
A quick corollary of the above lemma is that the smallest eigenvalue of $I-\text{diag}(\boldsymbol{\nu})^{-1}\Pi^{\top}\text{diag}(\boldsymbol{\mu})^{-1}\Pi$ is zero and is simple. As a result, we obtain an invertible matrix after dropping one of the rows.
@@
For any coupling matrix $\Pi$, the matrix $\text{diag}(\boldsymbol{\overline{\nu}})-\overline{\Pi}^{\top}\text{diag}(\boldsymbol{\mu})^{-1}\Pi$ is invertible, where $\boldsymbol{\overline{\nu}}$ and $\overline{\Pi}$ are the arrays after dropping the last rows.
The invertibility of the above matrix has been used in [36, 20, 6, 11] to remove the zero eigenvalues of $H(\Pi^{*})$ in the computation of the gradient of Sinkhorn distance (5). However, the ill-posedness in solving the linear systems is not addressed as the other eigenvalues are still close to 00. Next, we analyze the spectrum of the H-matrix, which guides the treatment of the ill-posedness.
@@ 4.2 Coupling matrices with uniform marginal distributions
Coupling matrices with uniform marginal distributions are of particular interest as they arise in many applications where the data are randomly sampled. They lead to H-matrices of the form
$H(\Pi):=\left[\begin{array}[]{ccccccccc}\text{diag}(\frac{1}{M}\mathbbm{1}_{M})&\Pi\\ (\Pi)^{\top}&\text{diag}(\frac{1}{N}\mathbbm{1}_{N})\end{array}\right]\in\mathbb{R}^{(M+N)\times(M+N)}.$ (48)
In this section, we compute the eigenvalues of the H-matrices for coupling matrices with uniform marginal distributions (including the permutation matrices) in Proposition 8. These eigenvalues shed light on the root of the ill-conditionedness of the H-matrices. In particular, if the coupling matrix is entrywise positive, we provide upper and lower bounds for the condition number of the H-matrix in terms of the spectral gap of $\Pi^{\top}\Pi$ in Theorem 9.
Proposition 8 (Eigenvalues of $H(\Pi)$ and singular values of $\Pi$).
Let $\Pi\in[0,1]^{M\times N}$ be a (not necessarily positive) coupling matrix with uniform marginal distributions. Let $M\leq N$ and assume $\Pi$ has rank $M$. Then, the eigenvalues (in descending order) and eigenvectors of $H$ defined in (48) are, for $j=1,\dots,M$,
$\lambda_{j}(H)$ $=\frac{\left(\frac{M+N}{MN}\right)+\sqrt{\left(\frac{M-N}{MN}\right)^{2}+4\sigma_{j}(\Pi)^{2}}}{2},\quad\mathbf{q}_{j}=\left[\begin{array}[]{ccccccccc}\frac{\kappa_{j}}{\sqrt{1+\kappa_{j}^{2}}}\mathbf{u}_{j}\\ \frac{1}{\sqrt{1+\kappa_{j}^{2}}}\mathbf{v}_{j}\end{array}\right],$ (49)
$\lambda_{M+1}(H)$ $=\dots=\lambda_{N}=\frac{1}{N},\quad\mathbf{q}_{i}=\left[\begin{array}[]{ccccccccc}\mathbb{0}\\ \boldsymbol{v}_{i}\end{array}\right],\,M<i\leq N,$
$\lambda_{N+M+1-j}(H)$ $=\frac{\left(\frac{M+N}{MN}\right)-\sqrt{\left(\frac{M-N}{MN}\right)^{2}+4\sigma_{j}^{2}(\Pi)}}{2},\quad\mathbf{q}_{N+M+1-j}=\left[\begin{array}[]{ccccccccc}\frac{-1}{\sqrt{1+\kappa_{j}^{2}}}\mathbf{u}_{j}\\ \frac{\kappa_{j}}{\sqrt{1+\kappa_{j}^{2}}}\mathbf{v}_{j}\end{array}\right],$
where $\sigma_{j}(\Pi)$ is the $j$-th singular value of $\Pi$ (in descending order) with $\mathbf{u}_{j}$ and $\mathbf{v}_{j}$ being the right and left singular vectors, $\{\mathbf{v}_{i}\}_{i=M+1}^{N}$ are the $N-M$ orthogonal vectors in the kernel of $\Pi$, and these scalars $\kappa_{j}=\left(\frac{\frac{1}{M}-\frac{1}{N}}{2\sigma_{j}(\Pi)}\right)+\sqrt{\left(\frac{\frac{1}{M}-\frac{1}{N}}{2\sigma_{j}(\Pi)}\right)^{2}+1}$.
In particular, when $M=N$ and $\Pi$ is a permutation coupling matrix, i.e., its rows are a permutation of the rows of $\frac{1}{N}I_{N}$. Then, the eigenvalues of $H(\Pi)$ are $\lambda_{1}(H)=\cdots=\lambda_{N}(H)=\frac{2}{N}$ and $\lambda_{N+1}(H)=\cdots=\lambda_{2N}(H)=0$.
@@
Denote an eigen-pair $H$ by $\left(\lambda,\left[\begin{array}[]{ccccccccc}\mathbf{u}\\ \mathbf{v}\end{array}\right]\right)$, where $\mathbf{u}\in\mathbb{R}^{M}$ and $\mathbf{v}\in\mathbb{R}^{N}$; that is, $\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbb{I}_{M}&\Pi\\ (\Pi)^{\top}&\frac{1}{N}\mathbb{I}_{N}\end{array}\right]\,\left[\begin{array}[]{ccccccccc}\mathbf{u}\\ \mathbf{v}\end{array}\right]=\lambda\left[\begin{array}[]{ccccccccc}\mathbf{u}\\ \mathbf{v}\end{array}\right]$.
Then, $\Pi\mathbf{v}=(\lambda-\frac{1}{M})\mathbf{u}$ and $\Pi^{\top}\mathbf{u}=(\lambda-\frac{1}{N})\mathbf{v}$. This implies that
$\left(\Pi(\Pi)^{\top}\right)\mathbf{u}=\left(\lambda-\frac{1}{M}\right)\left(\lambda-\frac{1}{N}\right)\mathbf{u},\quad\left((\Pi)^{\top}\Pi\right)\mathbf{v}=\left(\lambda-\frac{1}{M}\right)\left(\lambda-\frac{1}{N}\right)\mathbf{v}.$ (50)
That is, $\mathbf{u}$ and $\mathbf{v}$ are the right and left singular vectors of $\Pi$, corresponding to a singular value of $\Pi$ satisfying $\sigma_{j}^{2}=\left(\lambda-\frac{1}{M}\right)\left(\lambda-\frac{1}{N}\right)$, for some $1\leq j\leq M$. Solving this equation, we obtain two eigenvalues $\lambda_{\pm}=\frac{\left(\frac{1}{M}+\frac{1}{N}\right)\pm\sqrt{\left(\frac{1}{M}-\frac{1}{N}\right)^{2}+4\sigma_{j}^{2}}}{2}.$
To compute the eigenvector of $\lambda_{+}$, let $\mathbf{v}=\mathbf{v}_{j}$. Then $\Pi\mathbf{v}=\sigma_{j}\mathbf{u}_{j}=(\lambda_{+}-\frac{1}{M})\mathbf{u}$. So $\mathbf{u}=\frac{\sigma_{j}}{(\lambda_{+}-\frac{1}{M})}\mathbf{u}_{j}$ (recall that $\sigma_{j}\neq 0$ by $\text{rank}(\Pi)=m$). Note that $\frac{(\lambda_{+}-\frac{1}{N})}{\sigma_{j}}=\frac{\left(\frac{1}{M}-\frac{1}{N}\right)+\sqrt{\left(\frac{1}{M}-\frac{1}{N}\right)^{2}+4\sigma_{j}^{2}}}{2\sigma_{j}}$; hence, we have
$\frac{\sigma_{j}}{(\lambda_{+}-\frac{1}{M})}=\frac{(\lambda_{+}-\frac{1}{N})}{\sigma_{j}}=\frac{\frac{1}{M}-\frac{1}{N}}{2\sigma_{j}}+\sqrt{\left(\frac{\frac{1}{M}-\frac{1}{N}}{2\sigma_{j}}\right)^{2}+1}=\kappa_{j}$
@@
The above $\lambda_{\pm}$ account for $2M$ eigenvalues. The other $N-M$ eigenvalues correspond to $\sigma_{i}(\Pi)=0$, which has singular vectors $\mathbf{v}_{i}$, for $i=M+1,\dots,N$. Thus, setting $\mathbf{u}=\mathbb{0}$ and $\mathbf{v}=\mathbf{v}_{i}$, we have $\Pi^{\top}\mathbf{u}=\mathbb{0}=(\lambda-\frac{1}{N})\mathbf{v}_{i}$. So $\lambda=\frac{1}{N}$. This eigenvalue has multiplicity of $N-M$, with eigenvectors $\left[\begin{array}[]{ccccccccc}\mathbb{0}\\ \mathbf{v}_{i}\end{array}\right]$ for $i=M+1,\dots,N$.
At last, when $M=N$ and $\Pi$ is a permutation matrix, note that $\Pi^{\top}\Pi=\frac{1}{N^{2}}\mathbb{I}_{N}$. Thus, the singular values of $\Pi$ are $\sigma_{j}=\frac{1}{N}$ with multiplicity $N$. Applying (49), we obtain that the eigenvalues of $H$ are 00 and $\frac{2}{N}$, each with multiplicity $N$. ∎
Theorem 9 (Condition number of $H(\Pi)$).
Let $\Pi$ be a positive coupling matrix with uniform marginal distributions and singular values $\{\sigma_{k}\}$ in descending order. Then the condition number $\kappa(H)$ of $H$ in (48) is bounded by
$\frac{(M+N)^{2}}{2M^{2}N^{2}(\sigma_{1}^{2}-\sigma_{2}^{2})}\leq\kappa(H)\leq\frac{(M+N)^{2}}{M^{2}N^{2}(\sigma_{1}^{2}-\sigma_{2}^{2})}.$ (51)

### v2

@@ 4 Spectral analysis of the H-matrix
This section analyzes the spectrum of the matrix $\mathbf{H}$ in (2) and addresses the potential ill-posedness in solving the linear systems in (32), which are critical for computing the gradient and Hessian.
Recall that $\mathbf{H}$ is a block matrix consisting of the coupling matrix $\bm{\Pi}\in\mathbb{R}^{M\times N}$ and diagonal matrices of its two marginal vectors $\bm{\mu}$ and $\bm{\nu}$. We assume that the coupling matrix $\bm{\Pi}$ is known, either estimated via the Sinkhorn algorithm or computed analytically. We focus on two common types of coupling matrices: (i) a positive coupling matrix, where $\Pi_{ij}>0$ for all $i,j$, and (ii) a coupling matrix with uniform marginal distributions, where $\bm{\mu}=\mathbb{1}_{M}/M$ and $\bm{\nu}=\mathbb{1}_{N}/N$. The first type typically arises from the Sinkhorn algorithm, as all entries of the optimal coupling matrix are positive, as shown in (10) [15, 26]. The second type is common in applications with randomly sampled data. In entropic OT problems, the second type can be viewed as a special case of the first.
Our main contribution here is to establish upper and lower bounds for the condition number of $\mathbf{H}$ based on the spectral gap of $\bm{\Pi}^{\top}\bm{\Pi}$. We also provide examples to illustrate that $\mathbf{H}^{*}$ can be ill-conditioned. So the truncated SVD regularization, as discussed in Section 3.4, is necessary to mitigate numerical instability.
First, in Section 4.1, we prove in Theorem 6 that $\mathbf{H}$ is singular, with a simple zero eigenvalue, for any positive coupling matrix. Although removing one of the rows can eliminate the singularity, the linear system may still remain ill-posed.
Second, in Section 4.2, we analytically compute the eigenvalues of $\mathbf{H}$ for coupling matrices with uniform marginal distributions (including permutation matrices). In particular, we derive lower and upper bounds for the condition number of $\mathbf{H}$ in terms of the spectral gap of $\bm{\Pi}^{\top}\bm{\Pi}$, as established in Theorem 10. These bounds are extended to the $\mathbf{H}$-matrix computed by the Sinkhorn algorithm with early stopping in the entropic optimal transport (EOT) distance.
Finally, in Section 4.3, we derive the sharp asymptotic behavior of the condition number for one example in terms of the regularization strength $\epsilon$ and the sample size $N$, demonstrating that this linear system can be highly ill-conditioned. Interestingly, we observe similar asymptotic behavior empirically when applying to other randomly sampled datasets.
@@ 4.1 General positive coupling matrices
We show first that the $\mathbf{H}$ is singular with a simple zero eigenvalue for any positive coupling matrix $\bm{\Pi}$. As a result, we call $\kappa(\mathbf{H})=\frac{\lambda_{1}(\mathbf{H})}{\lambda_{N+M-1}(\mathbf{H})}$, the condition number of the matrix $\mathbf{H}$.
Theorem 6 (Simple zero eigenvalue for the $\mathbf{H}$-matrix).
For any positive coupling matrix $\bm{\Pi}\in\mathbb{R}^{M\times N}_{>0}$, the smallest eigenvalue of $\mathbf{H}$ is zero and it is simple, with eigenvector $\mathbf{q}_{0}=\left[\begin{array}[]{ccccccccc}\mathbb{1}_{M}\\ -\mathbb{1}_{N}\end{array}\right]$.
Its proof relies on the next Perron-Frobenius type lemma, which shows that the largest eigenvalue of the matrix $\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\,\text{diag}(\bm{\mu})^{-1}\bm{\Pi}$ is $1$ and is simple. We postpone the proof of the lemma to supplementary materials.
@@
For any coupling matrix with strictly positive entries $\bm{\Pi}\in\mathbb{R}^{M\times N}_{>0}$, the largest eigenvalue of matrix $\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\,\text{diag}(\bm{\mu})^{-1}\bm{\Pi}$ is $\lambda=1$ and has multiplicity one, with eigenvector $\mathbb{1}_{N}$. Similarly, the largest eigenvalue of the matrix $\text{diag}(\bm{\mu})^{-1}\bm{\Pi}\,\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}$ is $\lambda=1$ and has multiplicity one, with eigenvector $\mathbb{1}_{M}$.
@@
It is clear that 00 is an eigenvalue with eigenvector $\mathbf{q}_{0}$. We show next that any other eigenvector of 00 must be $\mathbf{q}_{0}$ up to a scalar factor. Note that the vector $\left[\begin{array}[]{ccccccccc}\bm{w}\\ \bm{v}\end{array}\right]$ is the eigenvector of 00 if and only if
$\left[\begin{array}[]{ccccccccc}\text{diag}(\bm{\mu})&\bm{\Pi}\\ \bm{\Pi}^{\top}&\text{diag}(\bm{\nu})\end{array}\right]\left[\begin{array}[]{ccccccccc}\bm{w}\\ \bm{v}\end{array}\right]=\mathbb{0}_{M+N}\Leftrightarrow\left[\begin{array}[]{ccccccccc}\text{diag}(\bm{\mu})\bm{w}+\bm{\Pi}\bm{v}=0\\ \text{diag}(\bm{\nu})\bm{v}+\bm{\Pi}^{\top}\bm{w}=0\end{array}\right],$ (43)
@@
$\bm{v}$ $=-\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\bm{w}=\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\text{diag}(\bm{\mu})^{-1}\bm{\Pi}\bm{v},$
$\bm{w}$ $=-\text{diag}(\bm{\mu})^{-1}\bm{\Pi}\bm{v}=\text{diag}(\bm{\mu})^{-1}\bm{\Pi}\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\bm{w}.$
By Lemma 7, $\bm{v}=a\mathbb{1}_{N}$ and $\bm{w}=b\mathbb{1}_{M}$ for some nonzero constant $a,b$ for each equation to hold. Plugging back to (43), we have $a=-b$, and $\left[\begin{array}[]{ccccccccc}\bm{w}\\ \bm{v}\end{array}\right]=a\mathbf{q}_{0}$. Thus, the zero eigenvalue of $\mathbf{H}$ is simple. ∎
A quick corollary of the above lemma is that the smallest eigenvalue of $\mathbb{I}-\text{diag}(\bm{\nu})^{-1}\bm{\Pi}^{\top}\text{diag}(\bm{\mu})^{-1}\bm{\Pi}$ is zero and is simple. As a result, we obtain an invertible matrix after dropping one of the rows.
@@
For any coupling matrix $\bm{\Pi}$, $\text{diag}(\bm{\overline{\nu}})-\overline{\bm{\Pi}}^{\top}\text{diag}(\bm{\mu})^{-1}\bm{\Pi}$ is invertible, where $\bm{\overline{\nu}}$ and $\overline{\bm{\Pi}}$ are the arrays after dropping the last rows.
The invertibility of the above matrix has been used in [38, 22, 7, 13] to remove the zero eigenvalues of $\mathbf{H}^{*}$ in the computation of the gradient of Sinkhorn distance (5). However, the ill-posedness in solving the linear systems is not addressed as the other eigenvalues are still close to 00. Next, we analyze the spectrum of the $\mathbf{H}$-matrix, which guides the treatment of the ill-posedness.
@@ 4.2 Coupling matrices with uniform marginal distributions
Coupling matrices with uniform marginal distributions are of particular interest as they arise in many applications. In this section, we compute the eigenvalues of the $\mathbf{H}$-matrices for coupling matrices $\bm{\Pi}$ with uniform marginal distributions (including the permutation matrices), i.e., $\bm{\Pi}\mathbb{1}=\frac{1}{M}\mathbb{1}_{M}$ and $\bm{\Pi}^{\top}\mathbb{1}=\frac{1}{N}\mathbb{1}_{N}$. These eigenvalues shed light on the root of the ill-conditionedness of the $\mathbf{H}$-matrices. In particular, if the coupling matrix is entrywise positive, we provide upper and lower bounds for the condition number of the $\mathbf{H}$-matrix in terms of the spectral gap of $\bm{\Pi}^{\top}\bm{\Pi}$ in Theorem 10.
Proposition 9 (Eigenvalues of $\mathbf{H}$ and singular values of $\bm{\Pi}$).
Let $\bm{\Pi}\in[0,1]^{M\times N}$ be a (not necessarily positive) coupling matrix with uniform marginal distributions. Let $M\leq N$ and assume $\bm{\Pi}$ has rank $M$. Then, the eigenvalues (in descending order) and eigenvectors of $\mathbf{H}$ defined in (2) are, for $j=1,\dots,M$,
$\lambda_{j}(\mathbf{H})$ $=\frac{\left(\frac{M+N}{MN}\right)+\sqrt{\left(\frac{M-N}{MN}\right)^{2}+4\sigma_{j}(\bm{\Pi})^{2}}}{2},\quad\mathbf{q}_{j}=\left[\begin{array}[]{ccccccccc}\frac{\kappa_{j}}{\sqrt{1+\kappa_{j}^{2}}}\mathbf{u}_{j}\\ \frac{1}{\sqrt{1+\kappa_{j}^{2}}}\mathbf{v}_{j}\end{array}\right],$ (44)
$\lambda_{M+1}(\mathbf{H})$ $=\dots=\lambda_{N}=\frac{1}{N},\quad\mathbf{q}_{i}=\left[\begin{array}[]{ccccccccc}\mathbb{0}\\ \mathbf{v}_{i}\end{array}\right],\,M<i\leq N,$
$\lambda_{N+M+1-j}(\mathbf{H})$ $=\frac{\left(\frac{M+N}{MN}\right)-\sqrt{\left(\frac{M-N}{MN}\right)^{2}+4\sigma_{j}^{2}(\bm{\Pi})}}{2},\quad\mathbf{q}_{N+M+1-j}=\left[\begin{array}[]{ccccccccc}\frac{-1}{\sqrt{1+\kappa_{j}^{2}}}\mathbf{u}_{j}\\ \frac{\kappa_{j}}{\sqrt{1+\kappa_{j}^{2}}}\mathbf{v}_{j}\end{array}\right],$
where $\sigma_{j}(\bm{\Pi})$ is the $j$-th singular value of $\bm{\Pi}$ (in descending order) with $\mathbf{u}_{j}$ and $\mathbf{v}_{j}$ being the right and left singular vectors, $\{\mathbf{v}_{i}\}_{i=M+1}^{N}$ are the $N-M$ orthogonal vectors in the kernel of $\bm{\Pi}$, and these scalars $\kappa_{j}=\left(\frac{\frac{1}{M}-\frac{1}{N}}{2\sigma_{j}(\bm{\Pi})}\right)+\sqrt{\left(\frac{\frac{1}{M}-\frac{1}{N}}{2\sigma_{j}(\bm{\Pi})}\right)^{2}+1}$.
In particular, when $M=N$ and $\bm{\Pi}$ is a permutation coupling matrix, i.e., its rows are a permutation of the rows of $\frac{1}{N}I_{N}$. Then, the eigenvalues of $H(\bm{\Pi})$ are $\lambda_{1}(H)=\cdots=\lambda_{N}(H)=\frac{2}{N}$ and $\lambda_{N+1}(H)=\cdots=\lambda_{2N}(H)=0$.
@@
Denote an eigen-pair $H$ by $\left(\lambda,\left[\begin{array}[]{ccccccccc}\mathbf{u}\\ \mathbf{v}\end{array}\right]\right)$, where $\mathbf{u}\in\mathbb{R}^{M}$ and $\mathbf{v}\in\mathbb{R}^{N}$; that is, $\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbb{I}_{M}&\bm{\Pi}\\ (\bm{\Pi})^{\top}&\frac{1}{N}\mathbb{I}_{N}\end{array}\right]\,\left[\begin{array}[]{ccccccccc}\mathbf{u}\\ \mathbf{v}\end{array}\right]=\lambda\left[\begin{array}[]{ccccccccc}\mathbf{u}\\ \mathbf{v}\end{array}\right]$.
Then, $\bm{\Pi}\mathbf{v}=(\lambda-\frac{1}{M})\mathbf{u}$ and $\bm{\Pi}^{\top}\mathbf{u}=(\lambda-\frac{1}{N})\mathbf{v}$. This implies that
$\left(\bm{\Pi}(\bm{\Pi})^{\top}\right)\mathbf{u}=\left(\lambda-\frac{1}{M}\right)\left(\lambda-\frac{1}{N}\right)\mathbf{u},\quad\left((\bm{\Pi})^{\top}\bm{\Pi}\right)\mathbf{v}=\left(\lambda-\frac{1}{M}\right)\left(\lambda-\frac{1}{N}\right)\mathbf{v}.$
That is, $\mathbf{u}$ and $\mathbf{v}$ are the right and left singular vectors of $\bm{\Pi}$, corresponding to a singular value of $\bm{\Pi}$ satisfying $\sigma_{j}^{2}=\left(\lambda-\frac{1}{M}\right)\left(\lambda-\frac{1}{N}\right)$, for some $1\leq j\leq M$. Solving this equation, we obtain two eigenvalues $\lambda_{\pm}=\frac{\left(\frac{1}{M}+\frac{1}{N}\right)\pm\sqrt{\left(\frac{1}{M}-\frac{1}{N}\right)^{2}+4\sigma_{j}^{2}}}{2}.$
To compute the eigenvector of $\lambda_{+}$, let $\mathbf{v}=\mathbf{v}_{j}$. Then $\bm{\Pi}\mathbf{v}=\sigma_{j}\mathbf{u}_{j}=(\lambda_{+}-\frac{1}{M})\mathbf{u}$. So $\mathbf{u}=\frac{\sigma_{j}}{(\lambda_{+}-\frac{1}{M})}\mathbf{u}_{j}$ (recall that $\sigma_{j}\neq 0$ by $\text{rank}(\bm{\Pi})=m$). Note that $\frac{(\lambda_{+}-\frac{1}{N})}{\sigma_{j}}=\frac{\left(\frac{1}{M}-\frac{1}{N}\right)+\sqrt{\left(\frac{1}{M}-\frac{1}{N}\right)^{2}+4\sigma_{j}^{2}}}{2\sigma_{j}}$; hence, we have
$\frac{\sigma_{j}}{(\lambda_{+}-\frac{1}{M})}=\frac{(\lambda_{+}-\frac{1}{N})}{\sigma_{j}}=\frac{\frac{1}{M}-\frac{1}{N}}{2\sigma_{j}}+\sqrt{\left(\frac{\frac{1}{M}-\frac{1}{N}}{2\sigma_{j}}\right)^{2}+1}=\kappa_{j}.$
@@
The above $\lambda_{\pm}$ account for $2M$ eigenvalues. The other $N-M$ eigenvalues correspond to $\sigma_{i}(\bm{\Pi})=0$, which has singular vectors $\mathbf{v}_{i}$, for $i=M+1,\dots,N$. Thus, setting $\mathbf{u}=\mathbb{0}$ and $\mathbf{v}=\mathbf{v}_{i}$, we have $\bm{\Pi}^{\top}\mathbf{u}=\mathbb{0}=(\lambda-\frac{1}{N})\mathbf{v}_{i}$. So $\lambda=\frac{1}{N}$. This eigenvalue has multiplicity of $N-M$, with eigenvectors $\left[\begin{array}[]{ccccccccc}\mathbb{0}\\ \mathbf{v}_{i}\end{array}\right]$ for $i=M+1,\dots,N$.
At last, when $M=N$ and $\bm{\Pi}$ is a permutation matrix, note that $\bm{\Pi}^{\top}\bm{\Pi}=\frac{1}{N^{2}}\mathbb{I}_{N}$. Thus, the singular values of $\bm{\Pi}$ are $\sigma_{j}=\frac{1}{N}$ with multiplicity $N$. Applying (44), we obtain that the eigenvalues of $H$ are 00 and $\frac{2}{N}$, each with multiplicity $N$. ∎
Theorem 10 (Condition number of $\mathbf{H}$).
Let $\bm{\Pi}$ be a positive coupling matrix with uniform marginal distributions and singular values $\{\sigma_{k}\}$ in descending order. Then the condition number $\kappa(\mathbf{H})$ of $\mathbf{H}$ in (2) is bounded by
$\frac{(M+N)^{2}}{2M^{2}N^{2}\rho(\bm{\Pi}^{\top}\bm{\Pi})}\leq\kappa(\mathbf{H})\leq\frac{(M+N)^{2}}{M^{2}N^{2}\rho(\bm{\Pi}^{\top}\bm{\Pi})},$ (45)
where $\rho(\bm{\Pi}^{\top}\bm{\Pi})$ is spectral gap of the matrix $\bm{\Pi}^{\top}\bm{\Pi}$.